OpenStudy (pottersheep):

Geometric Progressions Grade 11 Help please? I have to find the sum.

OpenStudy (pottersheep):

a = 1 , r = (p-1)/p and n = p Formula is : Sn = a(1 - r^n)/(1 - r) The answer is : (P^P - (p-1)^P) / P^P-1 i don't know how to get to the answer =/........can someone help me please

OpenStudy (pottersheep):

sorry, the answer is The answer is : (P^P - (p-1)^P) / P^(P-1), forgot a backet

OpenStudy (phi):

This is an algebra problem. \[\frac{1- \frac{(p-1)^p}{p^p}}{1- \frac{(p-1)}{p}}\]

OpenStudy (phi):

multiply top and bottom by p

OpenStudy (phi):

Do you see what happens to the bottom?

OpenStudy (anonymous):

@phi maybe p +1 instead of p? as the exponent i mean

OpenStudy (pottersheep):

hmm

OpenStudy (phi):

he said the exponent was n, and n was = p, unless I mis-read the question

OpenStudy (anonymous):

usually it mean you have n terms and sum is \[\frac{1-r^{n+1}}{1-r}\]

OpenStudy (phi):

Sn = a(1 - r^n)/(1 - r)

OpenStudy (anonymous):

oh my fault i did not read carefully. ignore me

OpenStudy (pottersheep):

n = p , and the formula we're given is a(1 - r^n)/(1 - r) :)

OpenStudy (pottersheep):

What happens to the denominator P^P if I multiply both top and bottom by p?

OpenStudy (phi):

did you try?

OpenStudy (anonymous):

although in my defense \[S_n=\frac{a(1-r^{n+1})}{1-r}\]

OpenStudy (pottersheep):

so the bottom part becomes -1/p

OpenStudy (phi):

The bottom is \[ 1- \frac{(p-1)}{p} \]

OpenStudy (pottersheep):

actually satellite73 Sn = a(1 - r^n)/(1 - r)

OpenStudy (pottersheep):

yeah

OpenStudy (phi):

so what is \[ p(1 - \frac{(p-1)}{p})\]

OpenStudy (pottersheep):

p - p - 1 ? = -1

OpenStudy (phi):

pretty close. you distribute the p (that means multiply every term in the parens by p to get p - (p-1) now distribute the - sign

OpenStudy (pottersheep):

opps! thanks for correcting me

OpenStudy (pottersheep):

I'll try to go on from here..thank you ~~

OpenStudy (phi):

so the bottom goes away (becomes 1), and dividing by 1 can be ignored. to finish, multiply the top by p then find a common denominator to simplify

OpenStudy (pottersheep):

YAY!!!!!!!!

OpenStudy (pottersheep):

Thank you!! :))